# Functional Regression with Nonstationarity and Error Contamination: Application to the Economic Impact of Climate Change

### v1

@@ A.2 Random Elements of Hilbert spaces
Appendix B Finite sample performance in a simulation study
@@ B.1 Simulation data generating process
We investigate the finite sample performance of the proposed estimator using the model (2) with generated nonstationary processes of $\{x_{t}\}_{t\geq 1}$ and $\{y_{t}\}_{t\geq 1}$. First, noting that $x_{t}$ can be written as
$x_{t}=\sum_{j=1}^{\infty}\langle x_{t},v_{j}\rangle v_{j}$ (B.27)
@@
for an orthonormal basis $\{v_{j}\}_{j=1}^{\infty}$ (to be specified later) of $\mathcal{H}$, and assuming that $\mathcal{H}^{N}=\operatorname{span}\{v_{1},\ldots,v_{d_{N}}\}$, we simulate realizations of $x_{t}$ by generating $\langle x_{t},v_{j}\rangle$ as a real-valued nonstationary (resp. stationary) process for each $j\leq d_{N}$ (resp. $j\geq d_{N}+1$). More specifically, we generate $\langle x_{t},v_{j}\rangle$ using the following AR(1) law of motion: for some $\alpha_{j}\neq 0$, $\beta_{j}\in(-1,1)$ and $\sigma_{\varepsilon,j}>0$,
@@
$\Delta\langle x_{t},v_{j}\rangle=\beta_{j}^{N}\Delta\langle x_{t-1},v_{j}\rangle+\sigma_{\varepsilon,j}\varepsilon_{j,t},\quad$ $j=1,\ldots,d_{N},$ (B.28)
$\langle x_{t},v_{j}\rangle=\alpha_{j}+\beta_{j}^{S}\langle x_{t-1},v_{j}\rangle+\sigma_{\varepsilon,j}\varepsilon_{j,t},\quad$ $j\geq d_{N}+1,$ (B.29)
@@
where $\varepsilon_{j,t}$ is iid $N(0,1)$ across $j$ and $t$, and also independent of any other variables. As will be detailed, $\sigma_{\varepsilon,j}$ is set to decay to zero as $j$ gets larger, and thus the time series $\langle x_{t},v_{j}\rangle$ in (B.29) has more importance in determining the properties of the stationary components of $x_{t}$ when $j$ is smaller. We first let $\beta_{j}^{N}$ be randomly determined in each simulation run, specifically as $\beta_{j}^{N}=s_{j}U_{j}^{N}$, where $U_{j}^{N}$ is a uniform random variable supported on [0.5,0.5][-0.5,0.5] (i.e., U[0.5,0.5]U[-0.5,0.5]), and $s_{j}$ is a Rademacher random variable independent of $U_{j}^{N}$; both sequences are independent across $j$. Moreover, given that (i) $\beta_{j}^{S}$ governs the correlation between $\langle{\rm P}^{S}x_{t},v_{j}\rangle$ and $\langle{\rm P}^{S}x_{t-\kappa},v_{j}\rangle$ and (ii) stationary time series tend to exhibit positive autocorrelation in many applications, we let $\beta_{j}^{S}$ be drawn independently from $U[0.4,0.9]$ for $j\leq M$, and from U[0.9,0.9]U[-0.9,0.9] for $j\geq M+1$, for some $M>0$ to be specified, in each repetition of the simulation experiment; combined with the decay of $\sigma_{\varepsilon,j}$, this ensures that the dominant part of the stationary components generally exhibits positive autocorrelation. The parameter $\sigma_{\varepsilon,j}$ determines the scale of $\langle x_{t},v_{j}\rangle$, which must decay to zero sufficiently fast for $C_{\kappa}^{S}$ to be a compact operator and hence well defined. We consider two simulation designs for this sequence, motivated by the setups in Seong and Seo, (2025). In the first design, referred to as the exponential design, we assume $\sigma_{\varepsilon,j}=1$ for $j\leq d_{N}+m$ and $\sigma_{\varepsilon,j}=(0.8)^{j-d_{N}-m}$ for $j=d_{N}+m+1,\ldots,d_{N}+M$, where $m$ (resp. $M$) is a moderately (resp. sufficiently) large integer. Given the required decay rate of the eigenvalues of $C_{\kappa}^{S}$ for our theoretical development, it is natural to consider the case where $\sigma_{\varepsilon,j}$ decreases geometrically for $j\geq M$; accordingly, we set $\sigma_{\varepsilon,j}=\sigma_{\varepsilon,M}(j-M)^{-2}$ for $j\geq M+1$. We use $m=7$ and $M=20$ throughout the simulation experiments. In the second design, referred to as the sparse design, we let $\sigma_{\varepsilon,j}=1$ for $j\leq d_{N}+m$, and $\sigma_{\varepsilon,j}=(0.1)^{j-d_{N}-m}$ for $j=d_{N}+m,\ldots,d_{N}+M$, with $M$ chosen to be sufficiently large. As in the exponential design, we set $\sigma_{\varepsilon,j}=\sigma_{\varepsilon,M}(j-M)^{-2}$ for $j\geq M+1$. It is expected that $(\widehat{D}_{\kappa})^{-1}_{{\mathsf{K}}}$ will tend to be more unstable under the sparse design, making it less favorable for the estimator. The intercepts $\alpha_{j}$ in (B.29) are independently generated from $N(0,1)$ in each simulation run. To generate $x_{t}$ as a function using (B.27) under these two simulation designs, we let $\{v_{j}\}_{j=1}^{\infty}$ be the Fourier basis functions, with the first eight basis functions randomly permuted in each repetition of the simulation.
@@
Similarly, noting that $y_{t}=\sum_{j=1}^{\infty}\langle y_{t},w_{j}\rangle w_{j}$ for an orthonormal basis $\{w_{j}\}_{j=1}^{\infty}$ of $\mathcal{H}_{y}$, we simulate $y_{t}$ by generating the coefficients $\langle y_{t},w_{j}\rangle$ and assigning a different set of Fourier basis functions to $\{w_{j}\}_{j=1}^{\infty}$, with a random permutation applied to the first eight functions in each simulation run. Throughout this simulation study, we assume that the linear map $f$ is defined by the following property: $fv_{j}=\gamma_{j}w_{j}$ for some $\gamma_{j}\neq 0$ for each $j$ but tending to zero as $j$ gets larger. It may be deduced from (2) that, in this case, $\langle y_{t},w_{j}\rangle=\gamma_{j}\langle x_{t},v_{j}\rangle+\langle u_{t},w_{j}\rangle$ for each $j$, and thus we generate $y_{t}$ as follows:
$y_{t}=\sum_{j=1}^{\infty}\langle y_{t},w_{j}\rangle w_{j},\quad\langle y_{t},w_{j}\rangle=\gamma_{j}\langle x_{t},v_{j}\rangle+\sigma_{u,j}u_{j,t},$
where $u_{j,t}$ is iid $N(0,1)$ across $j$ and $t$, and $\sigma_{u,j}$ is generated by the same mechanism as that of $\sigma_{\varepsilon,j}$. We let $\gamma_{j}=a_{j}U_{j}^{\gamma}$, where $U_{j}^{\gamma}$ is generated independently from U(1,1)U(-1,1), $a_{j}=1$ for $j\leq d_{N}$ and $a_{j}=(0.8)^{j-d_{N}}$ for $j=d_{N}+1$; the decay of $a_{j}$ is introduced to ensure the summability condition given in Assumption 3.
In computing the estimators, instead of $x_{t}$, we assume that we can only use $\tilde{x}_{t}=x_{t}+e_{t}$ with an additive measurement error $e_{t}$ given by $\sigma_{e}\eta_{t}$, where $\eta_{t}$ is the centered Brownian motion (as a function on $[0,1]$). The scalar $\sigma_{e}$ serves as a scale factor that controls the magnitude of the measurement error, and we let this depend on (the magnitude of) $x_{t}$. More specifically, we let $\sigma_{e}$ be chosen so that the nuclear norm of the covariance operator of $e_{t}$ matches $0\%$ (i.e., no measurement error), $5\%$ and $10\%$ of that of $\mathcal{E}_{t}^{x}=(\Delta{\rm P}^{N}x_{t},{\rm P}^{S}x_{t})$, which can be generated from the simulation DGP. Specifically, in each simulation run, the nuclear norm of the covariance operator of $\mathcal{E}_{t}^{x}$ (i.e., the sum of its eigenvalues) is approximated by the average of the corresponding sample estimates, computed from the simulated sequence of $\mathcal{E}_{t}^{x}$ based on (B.28) and (B.29). We use 400 repetitions to calculate the average in each simulation experiment. Naturally, a larger $\sigma_{e}$ corresponds to a larger measurement error.
@@ B.2 Simulation results
We examine the finite sample performance of our proposed estimators using the simulation DGP introduced in Section B.1. As in our empirical application, we consider the case where $d_{N}=2$, and compute our estimators with $\kappa=0$ and $\kappa=1$ to compare those in a few different scenarios on the magnitudes of the measuremenr errors. The tuning parameter $\mathsf{K}$} follows a pre-specified choice rule for the entire simulation experiments; specifically, given that ${\mathsf{K}}_{S}={\mathsf{K}}-d_{N}>0$ is required (note that this is a minimal requirement for nonzero $\widehat{f}_{\kappa}^{S}$ to be defined), we set ${\mathsf{K}}_{S}$ as ${\mathsf{K}}=d_{N}+\max_{j}\{\tilde{\lambda}_{j}>0.4\,T^{-0.2}\}$, where $\tilde{\lambda}_{j}$ is a scale-adjusted eigenvalue defined by $\tilde{\lambda}_{j}={\bm{\mathsf{\lambda}}_{j}[\widehat{D}_{\kappa}^{S}]}/{\sum_{j=1}^{\infty}\bm{\mathsf{\lambda}}_{j}[\widehat{D}_{\kappa}^{S}]}.$ [^4] Noting that any $\mathsf{K}$} satisfying Assumptions 2 and b necessarily depends on the scale of the functional observation $x_{t}$, the choice of $\mathsf{K}$} based on scale-adjusted eigenvalues was previously considered by Seong and Seo, (2025) as a scale-invariant selection in functional regression with stationary regressors. As a measure of the inaccuracy of the estimator $\widehat{f}_{\kappa}$ for $f$, we compute the Hilbert–Schmidt norm of $\widehat{f}_{\kappa}-f$, which can be calculated as $\sqrt{\sum_{j=1}^{\infty}\|\widehat{f}_{\kappa}(v_{j})-f(v_{j})\|^{2}}$ for any arbitrary orthonormal basis $\{v_{j}\}_{j=1}^{\infty}$ of $\mathcal{H}$. The simulation results are reported in Table 2. As may easily be expected from our theoretical results, the proposed estimator $\widehat{f}_{0}$ performs better than $\widehat{f}_{1}$ when there are no measurement errors. However, in the presence of measurement errors, $\widehat{f}_{0}$ not only performs worse than $\widehat{f}_{1}$ but also fails to show significant improvement as $T$ increases. Conversely, the performance of $\widehat{f}_{1}$ appears to be robust in the considered simulation setup, regardless of the presence of measurement errors. Overall, the simulation results support our theoretical findings in Section 4.
We have also examined the finite sample performance under a different set of parameters and obtained qualitatively similar results. As an example, we report the simulation results for the case where $d_{N}=3$ in Table 3.
@@
$\kappa=0$ 0.351 0.336 0.321 0.297 0.320 0.307 0.285 0.253
$\kappa=1$ 0.376 0.361 0.352 0.341 0.357 0.342 0.332 0.319
Magnitude of error: $5\%$
@@
$\kappa=0$ 0.438 0.432 0.434 0.429 0.430 0.426 0.428 0.423
$\kappa=1$ 0.411 0.375 0.354 0.340 0.400 0.360 0.337 0.321
Magnitude of error: $10\%$
@@
$\kappa=0$ 0.479 0.471 0.477 0.483 0.474 0.469 0.474 0.481
$\kappa=1$ 0.449 0.399 0.368 0.346 0.445 0.390 0.354 0.330
Notes: The average Hilbert Schmidt norm of $\widehat{f}-f$ is computed from 3000 Monte Carlo replications.
@@
$\kappa=0$ 0.338 0.321 0.310 0.287 0.303 0.290 0.275 0.246
$\kappa=1$ 0.372 0.349 0.343 0.332 0.349 0.330 0.324 0.309
Magnitude of error: $5\%$
@@
$\kappa=0$ 0.433 0.412 0.414 0.417 0.421 0.403 0.408 0.409
$\kappa=1$ 0.426 0.372 0.349 0.336 0.417 0.357 0.333 0.315
Magnitude of error: $10\%$
@@
$\kappa=0$ 0.479 0.455 0.460 0.474 0.471 0.449 0.456 0.472
$\kappa=1$ 0.478 0.403 0.364 0.341 0.477 0.396 0.353 0.324
Notes: The average Hilbert Schmidt norm of $\widehat{f}-f$ is computed from 3000 Monte Carlo replication.

### v3

@@ A.2 Random elements of Hilbert spaces
Appendix B Simulation study
@@ B.1 Simulation data generating process
We investigate the finite sample performance of the proposed estimator in simulation experiments. We assume that, if there were no measurement errors whatsoever, the observations would be given by $x_{t}=\mu_{x}+x_{t}^{0}$ and $y_{t}=\mu_{y}+y_{t}^{0}$, and the relationship $y_{t}^{0}=f(x_{t}^{0})+u_{t}$ holds, where $y_{t}^{0}$, $x_{t}^{0}$, and $u_{t}$ are generated processes whose unconditional means are zero. This leads to the model with an intercept $y_{t}=\mu+f(x_{t})+u_{t}$ with $\mu=\mu_{y}-f(\mu_{x})$ (see (24)). Noting that $x_{t}^{0}$ can be written as
$x_{t}^{0}=\sum_{j=1}^{\infty}\langle x_{t}^{0},v_{j}\rangle v_{j}$ (B.27)
@@
for an orthonormal basis $\{v_{j}\}_{j=1}^{\infty}$ (to be specified later) of $\mathcal{H}$ and assuming that $\mathcal{H}^{N}=\operatorname{span}\{v_{1},\ldots,v_{d_{N}}\}$, we simulate realizations of $x_{t}^{0}$ by generating $\langle x_{t}^{0},v_{j}\rangle$ as a real-valued nonstationary (resp. stationary) process for each $j\leq d_{N}$ (resp. $j\geq d_{N}+1$). Specifically, we generate $\langle x_{t}^{0},v_{j}\rangle$ using the following AR(1) law of motion: for some $\beta_{j}\in(-1,1)$ and $\sigma_{\varepsilon,j}>0$,
@@
$\Delta\langle x_{t}^{0},v_{j}\rangle=\beta_{j}^{N}\Delta\langle x_{t-1}^{0},v_{j}\rangle+\sigma_{\varepsilon,j}\varepsilon_{j,t},\quad$ $j=1,\ldots,d_{N},$ (B.28)
$\langle x_{t}^{0},v_{j}\rangle=\beta_{j}^{S}\langle x_{t-1}^{0},v_{j}\rangle+\sigma_{\varepsilon,j}\varepsilon_{j,t},\quad$ $j\geq d_{N}+1,$ (B.29)
@@
where $\varepsilon_{j,t}$ is iid $N(0,1)$ across $j$ and $t$, and also independent of any other variables. As will be detailed, $\sigma_{\varepsilon,j}$ is set to decay to zero as $j$ increases, and thus the time series $\langle x_{t}^{0},v_{j}\rangle$ in (B.29) becomes more important in determining the properties of the stationary components of $x_{t}^{0}$ when $j$ is smaller. We let $\beta_{j}^{N}$ be randomly determined in each simulation run, specifically as $\beta_{j}^{N}=s_{j}U_{j}^{N}$, where $U_{j}^{N}$ is a uniform random variable supported on [0.5,0.5][-0.5,0.5] (i.e., U[0.5,0.5]U[-0.5,0.5]), and $s_{j}$ is a Rademacher random variable independent of $U_{j}^{N}$; both sequences are independent across $j$. Moreover, given that (i) $\beta_{j}^{S}$ governs the correlation between $\langle{\rm P}^{S}x_{t}^{0},v_{j}\rangle$ and $\langle{\rm P}^{S}x_{t-\kappa}^{0},v_{j}\rangle$ and (ii) stationary time series tend to exhibit positive autocorrelation in many applications, we let $\beta_{j}^{S}$ be drawn independently from $U[0.5,0.9]$ for $j\leq M$, and from U[0.9,0.9]U[-0.9,0.9] for $j\geq M+1$, for some $M>0$ to be specified, in each repetition of the simulation experiment; combined with the decay of $\sigma_{\varepsilon,j}$, this ensures that the dominant part of the stationary components generally exhibits positive autocorrelation. The parameter $\sigma_{\varepsilon,j}$ determines the scale of $\langle x_{t}^{0},v_{j}\rangle$, which must decay to zero sufficiently fast for $C_{\kappa}^{S}$ to be a compact operator and hence well defined. We consider two simulation designs for this sequence, motivated by the setups in Seong and Seo, (2025). In the first design, referred to as the exponential design, we set $\sigma_{\varepsilon,j}=1$ for $j\leq d_{N}+m$ and $\sigma_{\varepsilon,j}=(0.8)^{j-d_{N}-m}$ for $j=d_{N}+m+1,\ldots,d_{N}+M$, where $m$ (resp. $M$) is a moderately (resp. sufficiently) large integer. Given the required decay rate of the eigenvalues of $C_{\kappa}^{S}$ for our theoretical development, it is natural to consider the case where $\sigma_{\varepsilon,j}$ decreases geometrically for $j\geq M$; accordingly, we set $\sigma_{\varepsilon,j}=\sigma_{\varepsilon,M}(j-M)^{-2}$ for $j\geq M+1$. We use $m=7$ and $M=20$ throughout the simulation experiments. In the second design, referred to as the sparse design, we set $\sigma_{\varepsilon,j}=1$ for $j\leq d_{N}+m$, and $\sigma_{\varepsilon,j}=(0.1)^{j-d_{N}-m}$ for $j=d_{N}+m+1,\ldots,d_{N}+M$. As in the exponential design, we set $\sigma_{\varepsilon,j}=\sigma_{\varepsilon,M}(j-M)^{-2}$ for $j\geq M+1$. To generate $x_{t}^{0}$ as a function using (B.27) under these two simulation designs, we use the Fourier basis functions, excluding the constant function, as in our empirical study in Section 5. We then construct $x_{t}=\mu_{x}+x_{t}^{0}$ by setting $\mu_{x}=\sum_{j=1}^{10}c_{x,j}v_{j}$, where $c_{x,j}$ is iid $N(0,1)$ across $j$ in each repetition of the simulation experiment.
@@
Similarly, we write $y_{t}^{0}=\sum_{j=1}^{\infty}\langle y_{t}^{0},w_{j}\rangle w_{j}$ for an orthonormal basis $\{w_{j}\}_{j=1}^{\infty}$ of $\mathcal{H}_{y}$. We simulate $y_{t}$ by generating the coefficients $\langle y_{t}^{0},w_{j}\rangle$ and using the Fourier basis functions, excluding the constant function, for $\{w_{j}\}_{j=1}^{\infty}$. Throughout this simulation study, we assume that the linear map $f$ is defined by the following property: $f(v_{j})=\gamma_{j}w_{j}$ for $\gamma_{j}\neq 0$ for each $j$, with $\gamma_{j}$ tending to zero as $j$ increases. We set $\gamma_{j}=a_{j}U_{j}^{\gamma}$, where $U_{j}^{\gamma}$ is generated independently from U(1,1)U(-1,1), $a_{j}=1$ for $j\leq d_{N}+m$ and $a_{j}=(0.8)^{j-d_{N}-m}$ for $j\geq d_{N}+m+1$; the decay of $a_{j}$ is introduced to ensure the summability condition in Assumption 3. Since $y_{t}^{0}=f(x_{t}^{0})+u_{t}$ holds, it may be deduced that, we in this case have $\langle y_{t}^{0},w_{j}\rangle=\gamma_{j}\langle x_{t}^{0},v_{j}\rangle+\langle u_{t},w_{j}\rangle$ for each $j$. We generate $y_{t}^{0}$ using this relationship and $u_{t}$ generated as follows:
$u_{t}$ $=\sum_{j=1}^{\infty}\langle u_{t},w_{j}\rangle w_{j},\quad\langle u_{t},w_{j}\rangle\sim_{iid}N(0,\sigma_{u,j}^{2}),$
where $\sigma_{u,j}$ is generated by the same mechanism as that of $\sigma_{\varepsilon,j}$. We then construct $y_{t}^{0}$ similarly as $x_{t}^{0}$, and then generate $y_{t}=\mu_{y}+y_{t}^{0}$ by setting $\mu_{y}=\sum_{j=1}^{10}c_{y,j}w_{j}$, where $c_{y,j}$ is iid $N(0,1)$ across $j$ in each simulation run.
Based on the above simulation DGP, constructed without any measurement errors, we consider the presence of measurement errors in $x_{t}$ and $y_{t}$, and hence assume that we can only observe $\tilde{x}_{t}=x_{t}+e_{t}$ and $\tilde{y}_{t}=y_{t}+e_{y,t}$ with additive measurement errors $e_{t}$ and $e_{y,t}$. In this case, from the equation $y_{t}=\mu+f(x_{t})+u_{t}$, we find that the following holds in the presence of measurement errors:
$\tilde{y}_{t}=\mu+f(\tilde{x}_{t})+\tilde{u}_{t},\quad\tilde{u}_{t}=u_{t}-f(e_{t})+e_{y,t}.$ (B.30)
The coefficient $f$ in the above equation is estimated using the generated observations $\{\tilde{x}_{t}\}_{t=1}^{T}$ and $\{\tilde{y}_{t}\}_{t=1}^{T}$ in each simulation run. We set $e_{t}=\sigma_{e}\eta_{t}$, where $\eta_{t}=\sum_{j=1}^{d_{N}+1}c_{\eta,j}v_{j}$ and $c_{\eta,j}$ is iid $N(0,1)$ across $j$. The scalar $\sigma_{e}$ serves as a scale factor that controls the magnitude of the measurement error, and we let it depend on (the magnitude of) $\mathcal{E}_{t}^{x}$. Specifically, we choose $\sigma_{e}$ so that the nuclear norm of the covariance operator of $e_{t}$ matches $0\%$ (i.e., no measurement errors in $x_{t}$), $50\%$ and $100\%$ of that of $\mathcal{E}_{t}^{x}=(\Delta{\rm P}^{N}x_{t},{\rm P}^{S}x_{t})$, which can be generated from the simulation DGP. In each simulation run, the nuclear norm of the covariance operator of $\mathcal{E}_{t}^{x}$ (i.e., the sum of its eigenvalues) is approximated by the average of the corresponding sample estimates, computed from the simulated sequence of $\mathcal{E}_{t}^{x}$ based on (B.28) and (B.29); we use 800 repetitions to calculate the average. We similarly set $e_{y,t}=\sigma_{e,y}\xi_{t}$, where $\xi_{t}=\sum_{j=1}^{d_{N}+1}c_{\xi,j}w_{j}$ and $c_{\xi,j}$ is iid $N(0,1)$ across $j$. However, since $e_{t}$ and $u_{t}$ are independent in the simulation setup, a larger measurement error in $x_{t}$ unintentionally tends to be accompanied by a larger magnitude (variance) of $\tilde{u}_{t}=u_{t}-f(e_{t})+e_{y,t}$, which in turn may affect the finite sample performance of the considered estimator and statistical inference. We thus choose $\sigma_{e,y}$ so that the nuclear norm of the covariance operator of $e_{y,t}$ matches $(100-\alpha)\%$ of that of $\mathcal{E}_{t}^{x}=(\Delta{\rm P}^{N}x_{t},{\rm P}^{S}x_{t})$ when $\sigma_{e}$ matches $\alpha\%$ of it for $e_{t}$. This approach allows us to control significant changes in the magnitude of the error $\tilde{u}_{t}$ in (B.30). Of course, unlike $e_{t}$, $e_{y,t}$ alone does not cause problematic issues, such as inconsistency, in the standard covariance-based estimator $\widehat{f}_{0}$. Thus, the case with $\sigma_{e}=0$ corresponds to the error-free scenario, which we consider in Assumption E.
@@ B.2 Simulation results
We examine the finite sample performance of our proposed estimators using the simulation DGP introduced in Section B.1. As in our empirical application, we consider the case where $d_{N}=2$, and compute our estimators with $\kappa=0$ and $\kappa=1$ to compare them in a few different scenarios regarding the scales of measurement errors. The tuning parameter $\mathsf{K}$} follows a pre-specified choice rule for the entire simulation experiments; specifically, given that ${\mathsf{K}}_{S}={\mathsf{K}}-d_{N}>0$ is required (note that this is a minimal requirement for nonzero $\widehat{f}_{\kappa}^{S}$ to be defined), we set $\mathsf{K}$} as ${\mathsf{K}}=d_{N}+\max_{j}\{\tilde{\lambda}_{j}>0.4\,T^{-0.2}\}$, where $\tilde{\lambda}_{j}$ is a scale-adjusted eigenvalue defined by $\tilde{\lambda}_{j}={\bm{\mathsf{\lambda}}_{j}[\widehat{D}_{\kappa}^{S}]}/{\sum_{j=1}^{\infty}\bm{\mathsf{\lambda}}_{j}[\widehat{D}_{\kappa}^{S}]}.$ [^4] Noting that any $\mathsf{K}$} satisfying Assumptions 2 and b necessarily depends on the scale of the functional observation $x_{t}$, the choice of $\mathsf{K}$} based on scale-adjusted eigenvalues was previously considered by Seong and Seo, (2025) as a scale-invariant selection in functional regression with stationary regressors. As a measure of the inaccuracy of the estimator $\widehat{f}_{\kappa}$ for $f$, we compute the Hilbert–Schmidt norm of $\widehat{f}_{\kappa}-f$, which can be calculated as $\sqrt{\sum_{j=1}^{\infty}\|\widehat{f}_{\kappa}(v_{j})-f(v_{j})\|^{2}}$ for any arbitrary orthonormal basis $\{v_{j}\}_{j=1}^{\infty}$ of $\mathcal{H}$. The simulation results are reported in Table 2. As may easily be expected from our theoretical results, the proposed estimator $\widehat{f}_{0}$ performs better than $\widehat{f}_{1}$ when there are no measurement errors. However, in the presence of measurement errors, $\widehat{f}_{0}$ performs worse than $\widehat{f}_{1}$, and the performance gap increases with the scale of measurement errors. The performance of $\widehat{f}_{1}$ appears to be robust in the considered simulation setup, regardless of the scale of measurement errors. Overall, the simulation results support our theoretical findings in Section 4.
In addition to evaluating the overall accuracy of the estimators, we also examine the coverage probability of the 95% confidence interval for $\langle f(\zeta),\varphi\rangle$, based on the pointwise asymptotic normality result (21), which follows from Theorems 4.2 and 4.4 (and their extension discussed in Section 4.4). We let $\zeta=\sum_{j=1}^{\infty}c_{\zeta,j}v_{j}$, where $c_{\zeta,j}=1$ for $j=1,\ldots,9$, and $c_{\zeta,j}=(j-8)^{-2}$ thereafter, and let $\varphi$ be $w_{1}$. The results are reported in Table 3. As expected, $\widehat{f}_{0}$ yields poor coverage probabilities in the presence of measurement errors in $x_{t}$, while $\widehat{f}_{1}$ provides improved results, and its performance is robust across different scales of measurement errors in $x_{t}$.
We also examined the finite sample performance under a different set of parameters and obtained qualitatively similar results. As an example, we report the simulation results for the case where $d_{N}=3$ in Tables 4 and 5.
@@
$\kappa=0$ 0.957 0.918 0.879 0.830 0.941 0.904 0.863 0.807
$\kappa=1$ 1.019 0.966 0.932 0.901 1.021 0.971 0.937 0.910
Scale of measurement errors: $50\%$
@@
$\kappa=0$ 1.072 1.051 1.029 1.005 1.031 1.001 0.972 0.940
$\kappa=1$ 1.010 0.957 0.926 0.897 1.010 0.964 0.937 0.907
Scale of measurement errors: $100\%$
@@
$\kappa=0$ 1.184 1.179 1.171 1.154 1.135 1.120 1.100 1.076
$\kappa=1$ 1.001 0.948 0.920 0.894 0.990 0.956 0.933 0.905
Notes: The average Hilbert Schmidt norm of $\widehat{f}-f$ is computed from 3000 Monte Carlo replications. The scale of measurement errors is given by $\alpha\%$ when $\sigma_{e}$ is chosen so that the nuclear norm of the covariance operator of $e_{t}$ matches $\alpha\%$ of that of $\mathcal{E}_{t}^{x}=(\Delta{\rm P}^{N}x_{t},{\rm P}^{S}x_{t})$.
Table 3: The coverage probability of the 95% confidence interval for $\langle f(\zeta),\varphi\rangle$ when $d_{N}=2$.
Exponential design Sparse design
Scale of measurement errors: $0\%$
Estimators $T=100$ $200$ $400$ $800$ $T=100$ $200$ $400$ $800$
$\kappa=0$ 0.768 0.854 0.917 0.934 0.798 0.868 0.918 0.940
$\kappa=1$ 0.774 0.854 0.917 0.930 0.805 0.867 0.922 0.935
Scale of measurement errors: $50\%$
Estimators $T=100$ $200$ $400$ $800$ $T=100$ $200$ $400$ $800$
$\kappa=0$ 0.615 0.689 0.763 0.838 0.677 0.765 0.831 0.883
$\kappa=1$ 0.738 0.843 0.914 0.928 0.779 0.864 0.927 0.933
Scale of measurement errors: $100\%$
Estimators $T=100$ $200$ $400$ $800$ $T=100$ $200$ $400$ $800$
$\kappa=0$ 0.158 0.222 0.307 0.404 0.238 0.335 0.440 0.576
$\kappa=1$ 0.582 0.781 0.892 0.928 0.668 0.818 0.892 0.937
Notes: The coverage probability is computed from 3000 Monte Carlo replications. The scale of measurement errors is determined as described in Table 2.
@@
$\kappa=0$ 0.936 0.896 0.863 0.819 0.916 0.878 0.848 0.797
$\kappa=1$ 1.021 0.951 0.919 0.894 1.026 0.954 0.929 0.902
Scale of measurement errors: $50\%$
@@
$\kappa=0$ 1.035 1.009 0.987 0.957 0.995 0.969 0.938 0.903
$\kappa=1$ 1.005 0.943 0.915 0.891 1.007 0.946 0.925 0.900
Scale of measurement errors: $100\%$
@@
$\kappa=0$ 1.143 1.135 1.122 1.100 1.096 1.080 1.058 1.030
$\kappa=1$ 0.989 0.931 0.909 0.887 0.981 0.936 0.920 0.899
Notes: The average Hilbert Schmidt norm of $\widehat{f}_{\kappa}-f$ is computed from 3000 Monte Carlo replication. The scale of measurement errors is determined as described in Table 2.
Table 5: The coverage probability of the 95% confidence interval for $\langle f(\zeta),\varphi\rangle$ when $d_{N}=3$.
Exponential design Sparse design
Scale of measurement errors: $0\%$
Estimators $T=100$ $200$ $400$ $800$ $T=100$ $200$ $400$ $800$
$\kappa=0$ 0.733 0.808 0.884 0.926 0.767 0.838 0.906 0.924
$\kappa=1$ 0.735 0.811 0.887 0.928 0.772 0.840 0.907 0.931
Scale of measurement errors: $50\%$
Estimators $T=100$ $200$ $400$ $800$ $T=100$ $200$ $400$ $800$
$\kappa=0$ 0.577 0.657 0.740 0.806 0.641 0.716 0.794 0.866
$\kappa=1$ 0.700 0.805 0.878 0.918 0.751 0.830 0.896 0.928
Scale of measurement errors: $100\%$
Estimators $T=100$ $200$ $400$ $800$ $T=100$ $200$ $400$ $800$
$\kappa=0$ 0.162 0.224 0.305 0.435 0.227 0.320 0.429 0.572
$\kappa=1$ 0.517 0.712 0.842 0.914 0.632 0.762 0.872 0.928
Notes: The coverage probability is computed from 3000 Monte Carlo replications. The scale of measurement errors is determined as described in Table 2.